# Comfortability of quantum walks on embedded graphs on surfaces

### v1

@@ 7 Comfortability (Proof of main theorem)
In this section, we subject to Assumption 1. Note that because of the hedgehog boundary condition, the set of tails has a bijection map to the set of bridges. Then for a bridge arc $e\in A_{br}$, put $t^{2}(e):=t(\mathrm{is}^{\sharp}(\bar{e}))\in\delta X$. Let us set $\eta\in\mathbb{C}^{A_{br}}$ by
$\eta(e):=\frac{1}{bc\omega}(Q\boldsymbol{\alpha}_{in})(t^{2}(e)).$
Here the matrix $Q:=S-dI$ represents the scattering after a quantum walker penetrates the interior at least once. By using the notation $\eta$ of such a scattering, the stationary state $\psi_{\infty}$ is expressed as follows.
Lemma 7.1.
Under Assumption 1, for any bridge arc $e\in A_{br}$,
$\psi_{\infty}(e)=(-1)^{\tau(e)}\;\omega\;(\;\eta(e)+d\eta(\bar{e})\;).$
@@
$f:=(\;\xi_{0}^{+},e_{0},\;\;\xi_{1}^{-},\xi_{1}^{+},e_{1},\;\;\xi_{2}^{-},\xi_{2}^{+},e_{2},\;\dots\\ \dots,\;\xi_{\kappa-1}^{-},\xi_{\kappa-1}^{-},e_{\kappa-1},\;\;\xi_{0}^{-}\;),$
the stationary state at the island arc is
@@
$\psi_{\infty}(\xi_{m+1}^{-})=(1/c)(S_{f}-dI_{f})\boldsymbol{\alpha}_{f}(m+1)=\omega b\eta(e_{m}).$ (7.27)
by Theorem 1.1. Lemma 6.1 leads
$\psi_{\infty}(\xi_{m}^{+})=(-1)^{\tau(e_{m})}\omega^{-1}\psi_{\infty}(\xi_{m+1}^{-})=(-1)^{\tau(e_{m})}b\eta(e_{m})$ (7.28)
By (6.23), under the assumption of $d\in\mathbb{R}$, we have
$\begin{bmatrix}\psi(e_{1}^{R})\\ \psi(e_{1}^{L})\end{bmatrix}$ $=\frac{1}{b}\begin{bmatrix}d^{\prime}&-b\\ 1&0\end{bmatrix}\begin{bmatrix}-a&1\\ -c^{\prime}&0\end{bmatrix}\begin{bmatrix}\psi(e_{0}^{R})\\ \psi(e_{0}^{L})\end{bmatrix}$
$=\frac{1}{b}\begin{bmatrix}1&d^{\prime}\\ d^{\prime}&1\end{bmatrix}\begin{bmatrix}\omega^{\prime}&0\\ 0&1\end{bmatrix}\begin{bmatrix}\psi(e_{0}^{R})\\ \psi(e_{0}^{L}),\end{bmatrix}$
which implies
$\psi(e_{1}^{R})=\frac{\omega^{\prime}}{b}(\psi(e_{0}^{R})+d^{\prime}\psi(e_{2}^{L})),\;\psi(e_{1}^{L})=\frac{\omega^{\prime}}{b}(\psi(e_{0}^{R})+d^{\prime}\psi(e_{2}^{L}))$ (7.29)
by Lemma 6.1. Recalling that the arcs $e_{1}^{R}$, $e_{1}^{L}$ are bridge arcs with $\bar{e_{1}^{L}}=e_{1}^{R}=e$ and $e_{0}^{R}=\mathrm{is}(e)$, $e_{0}^{L}=\mathrm{is}^{\sharp}(e)$, we obtain
$\psi_{\infty}(e)=(-1)^{\tau(e)}\omega(\eta(e)+\eta(\bar{e}))$
by inserting (7.27) and (7.28) into (7.29). ∎
Let $\sigma:\mathbb{C}^{A_{br}}\to\mathbb{C}^{A_{br}}$ be the flip-flop matrix such that
$(\sigma\psi)(a)=\psi(\bar{a})$
@@
Here the scattering matrix $S$ is regarded as the operator on $\mathbb{C}^{A_{br}}$ with the bijection map $A_{br}\to\delta X$ by $x=t(\mathrm{is}^{\sharp}(\bar{e}))\in\delta X$ for any $e\in A_{br}$(: note that the boundary is the hedgehog). By using the expression of the stationary state $\psi_{\infty}$ in Lemma 7.1, the comfortability $\mathcal{E}$ is described as follows.
Lemma 7.2.
@@
$\mathcal{E}=\frac{1}{|c|^{2}}||Q\boldsymbol{\alpha}||^{2}+\frac{1}{2|bc|^{2}}||(\sigma+dI)Q\boldsymbol{\alpha}||^{2}.$
@@
    Island: Since every island arc belongs to unique facial walk, $\mathcal{E}^{island}$ can be further decomposed into each facial walk by
@@
    where $\mathcal{E}_{f}^{island}=\sum_{e\in f\cap A_{is}}|\psi_{\infty}(e)|^{2}$. By Lemma 7.1, the comfortability on the island is deformed by
    $\mathcal{E}_{f}^{island}$ $=\frac{1}{2}\sum_{m=0}^{|f|-1}\left\{\;|(-1)^{\tau(e_{m})}b\eta_{m}|^{2}+|\omega b\eta_{m}|^{2}\;\right\}$
@@
    Bridge: By Lemma 7.1, we have
    $\mathcal{E}^{bridge}$ $=\frac{1}{2}\sum_{e\in A_{br}}|(-1)^{\tau(e)}\omega(\eta(e)+d\eta(\bar{e}))|^{2}$
    $=\frac{1+d^{2}}{2}\sum_{e\in A_{br}}|\eta(e)|^{2}+d\cdot\sum_{e\in A_{br}}\mathrm{Re}\left[\sum_{e\in A_{br}}\eta(e)\overline{\eta(\bar{e})}\right]$
    $=\frac{1}{2}\left\{(1+d^{2})\;||\eta||^{2}+2d\cdot\langle\eta,\sigma\eta\rangle\right\}$
@@
Now let us set the inflow inserting $1$ from a tail which is chosen uniformly at random, that is, an inflow $\boldsymbol{\alpha}_{in}$ is selected randomly from
$\{\delta_{e}\;|\;e\in A^{br}\}.$
@@
Under the Assumption 1 and the uniformly at random inflow, the average of the comfortability is expressed by
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(\overline{a\omega})^{\mathrm{dist}_{f}(\bar{e},e)}\;(1-|a|^{2\mathrm{dist}_{f}(e,\bar{e})})\;\bigg{\}}.$
Here $|f|$ is the length of the facial walk $f$ in $(G,\rho,\tau)$ and $\mathrm{dist}_{f}(e,e^{\prime})$ is the distance from $e$ to $e^{\prime}$ along the facial walk $f$ in $(G,\rho,\tau)$ for any $e,e^{\prime}\in f$.
@@
By Lemma 7.2, it holds that
$\mathbb{E}[\mathcal{E}]$ $=\frac{1}{|c|^{2}}\mathbb{E}[\;||Q\boldsymbol{\alpha}||^{2}\;]+\frac{1}{|bc|^{2}}\mathbb{E}[\;||(\sigma+dI)Q\boldsymbol{\alpha}||^{2}\;]$
@@
Now to describe the above RHS more explicitly, let us compute $QQ^{*}$ as follows. Note that $Q=\oplus_{f\in F}Q_{f}$, and for each $f\in F$ with $f=(e_{0},e_{1},\dots,{e_{|f|-1}})$, $Q_{f}$ can be expanded by
@@
since $P_{f}^{|f|}=I_{f}$. Then we have
$Q_{f}Q_{f}^{*}$ $=\frac{|bc|^{2}}{|1-(a\omega)^{n}|^{2}}\sum_{k=0}^{n-1}\left(\sum_{{\scriptsize\begin{matrix}\ell,m\in\mathbb{Z}_{n}\\ \text{ with }m-\ell=k\end{matrix}}}(a\omega)^{m}(\overline{a\omega})^{\ell}\right)P_{f}^{k}.$ (7.33)
Here $w_{k}:=\sum_{\ell,m\in\mathbb{Z}_{n},\;m-\ell=k}\;(a\omega)^{m}(\overline{a\omega})^{\ell}$ is reduced to
$w_{k}=\frac{1}{|b|^{2}}\left\{(a\omega)^{k}(1-|a|^{2(|f|-k)})+(\overline{a\omega})^{|f|-k}(1-|a|^{2k})\right\}.$
Inserting the above expression for $w_{k}$ into (7.33), we obtain
$(Q_{f}Q_{f}^{*})_{e_{\ell},e_{m}}=(-1)^{\int_{e_{m}}^{e_{\ell}}\tau}\frac{|b|^{2}}{|1-(a\omega)^{|f|}|^{2}}\\ \times\left\{(a\omega)^{m-\ell}(1-|a|^{2(|f|-(m-\ell))})+(\overline{a\omega})^{|f|-(m-\ell)}(1-|a|^{2(m-\ell)})\right\}.$ (7.34)
By using (7.34), $\mathrm{tr}(QQ^{*})$ and $\mathrm{tr}(QQ^{*}\sigma)$ are described by
@@
and
$\mathrm{tr}(QQ^{*}\sigma)$ $=\sum_{e}\sum_{e^{\prime}}(QQ^{*})_{e,e^{\prime}}(\sigma)_{e^{\prime},e}=\sum_{e,\bar{e}}(QQ^{*})_{\bar{e},e}$
$=\sum_{f\in F}\sum_{e\in f\cap\bar{f}}(QQ^{*})_{\bar{e},e}$
@@
$\qquad\sum_{e\in f\cap\bar{f}}\left\{(a\omega)^{\mathrm{dist}_{f}(e,\bar{e})}(1-|a|^{2\mathrm{dist}_{f}(\bar{e},e)})+(a\omega)^{\mathrm{dist}_{f}(\bar{e},e)(1-|a|^{2\mathrm{dist}(e,\bar{e})})}\right\}$ (7.36)
Here the last equality derives from $\mathrm{dist}_{f}(\bar{e},e)=|f|-\mathrm{dist}_{f}(e,\bar{e})$ and $(-1)^{\int_{e}^{\bar{e}}}\tau=1$ since $\bar{e}\in F$ and $e,\bar{e}$ are in the same sheet for any $e\in f\cap\bar{f}$. Combining (7) with (7.35) and (7.36), we obtain the desired conclusion. ∎

### v2

@@ 7 Comfortability (Proof of main theorem)
In this section, we discuss the comfortability under Assumption 1. Note that because of the hedgehog boundary condition, the set of tails has a bijection map to the set of bridges $\varphi$ in (6.36). Let us set $\eta\in\mathbb{C}^{A_{br}}$ by
$\eta(e):=\frac{1}{bc\omega}(Q\boldsymbol{\alpha}_{in})(t(\mathrm{is}^{\sharp}(\bar{e}))).$
See (4.15) and Figs. 9, 11 for the difinitions of $\mathrm{is}(e)$ and $\mathrm{is}^{\sharp}(e)$. Here the matrix $Q:=S-dI$ represents the scattering after a quantum walker penetrates the interior at least once. By using the notation $\eta$ of such a scattering, the stationary state $\psi_{\infty}$ is expressed as follows.
Proposition 7.1.
Under Assumption 1, for any bridge arc $e\in A_{br}$, we have
$\psi_{\infty}(e)=\;\omega\;(\;\eta(e)+(-1)^{\tau(e)}d\eta(\bar{e})\;).$
@@
$f:=(\;\xi_{0}^{+},e_{0},\;\;\xi_{1}^{-},\xi_{1}^{+},e_{1},\;\;\xi_{2}^{-},\xi_{2}^{+},e_{2},\;\dots\\ \dots,\;\xi_{\kappa-1}^{-},\xi_{\kappa-1}^{+},e_{\kappa-1},\;\;\xi_{0}^{-}\;),$
where $(\xi_{j}^{-},\xi_{j}^{+})\in\delta A$ and $e_{j}\in A_{br}$, the stationary state at the island arc is
@@
$\psi_{\infty}(\xi_{m+1}^{-})=(1/c)(S_{f}-dI_{f})\boldsymbol{\alpha}_{in}(m+1)=\omega b\eta(e_{m}).$ (7.37)
by the definition of the scattering matrix in (6.25). Lemma 6.1 leads
$\psi_{\infty}(\xi_{m}^{+})=(-1)^{\tau(e_{m})}\omega^{-1}\psi_{\infty}(\xi_{m+1}^{-})=(-1)^{\tau(e_{m})}b\eta(e_{m}).$ (7.38)
Under the assumption of $d\in\mathbb{R}$, we have
$\begin{bmatrix}\psi_{\infty}(e_{1}^{R})\\ \psi_{\infty}(e_{1}^{L})\end{bmatrix}$ $=\frac{1}{b}\begin{bmatrix}d^{\prime}&-b\\ 1&0\end{bmatrix}\begin{bmatrix}-a&1\\ -c^{\prime}&0\end{bmatrix}\begin{bmatrix}\psi_{\infty}(e_{0}^{R})\\ \psi_{\infty}(e_{0}^{L})\end{bmatrix}$
$=\frac{1}{b}\begin{bmatrix}1&d^{\prime}\\ d^{\prime}&1\end{bmatrix}\begin{bmatrix}\omega^{\prime}&0\\ 0&1\end{bmatrix}\begin{bmatrix}\psi_{\infty}(e_{0}^{R})\\ \psi_{\infty}(e_{0}^{L})\end{bmatrix}$
$=\frac{\omega^{\prime}}{b}\begin{bmatrix}1&d^{\prime}\\ d^{\prime}&1\end{bmatrix}\begin{bmatrix}\psi_{\infty}(e_{0}^{R})\\ \psi_{\infty}(e_{2}^{L})\end{bmatrix}.$
Here we used (6.29), (6.30) and (6.28) in the first, the second and the third equalities, respectively. Then it holds that
$\psi_{\infty}(e_{1}^{R})=\frac{\omega^{\prime}}{b}(\psi_{\infty}(e_{0}^{R})+d^{\prime}\psi_{\infty}(e_{2}^{L})),\;\psi_{\infty}(e_{1}^{L})=\frac{\omega^{\prime}}{b}(d^{\prime}\psi_{\infty}(e_{0}^{R})+\psi_{\infty}(e_{2}^{L})).$ (7.39)
Remark that if we put $e_{1}^{R}=e_{m}\in A_{br}$, then
$\psi_{\infty}(e_{0}^{R})=(-1)^{\tau(e_{m})}b\eta(e_{m})\text{ and}\psi_{\infty}(e_{2}^{L})=(-1)^{\tau(\bar{e}_{m})}b\eta(\bar{e}_{m})$
hold by (7.38). Then we obtain
$\psi_{\infty}(e_{m})=\omega(\eta(e_{m})+(-1)^{\tau(e_{m})}d\eta(\bar{e}_{m}))$
by inserting (7.38) into (7.39), which is the desired conclusion. ∎
Let $\sigma:\mathbb{C}^{A_{br}}\to\mathbb{C}^{A_{br}}$ be the flip-flop with the twist $\tau$ such that
$(\sigma\psi)(a)=(-1)^{\tau(a)}\psi(\bar{a})$
@@
Here the scattering matrix $S$ is regarded as the operator on $\mathbb{C}^{A_{br}}$ with the bijection map $A_{br}\to\delta X$ by $x=t(\mathrm{is}^{\sharp}(\bar{e}))\in\delta X$ for any $e\in A_{br}$(: note that the boundary is the hedgehog). By using the expression of the stationary state $\psi_{\infty}$ in Proposition 7.1, the comfortability $\mathcal{E}$ is described as follows.
Proposition 7.2.
@@
$\mathcal{E}=\frac{1}{|c|^{2}}||(S-dI)\boldsymbol{\alpha}_{in}||^{2}+\frac{1}{2|bc|^{2}}||(\sigma+dI)(S-dI)\boldsymbol{\alpha}_{in}||^{2}.$
@@
    Island: Since every island arc belongs to unique facial walk, $\mathcal{E}^{island}$ can be further decomposed into facial walks by
@@
    where $\mathcal{E}_{f}^{island}=\sum_{e\in f\cap A_{is}}|\psi_{\infty}(e)|^{2}$. By Proposition 7.1, the comfortability on the island is deformed by
    $\mathcal{E}_{f}^{island}$ $=\frac{1}{2}\sum_{m=0}^{|f|-1}\left\{\;|(-1)^{\tau(e_{m})}b\eta(e_{m})|^{2}+|\omega b\eta(e_{m})|^{2}\;\right\}$
@@
    Bridge: By Proposition 7.1, we have
    $\mathcal{E}^{bridge}$ $=\frac{1}{2}\sum_{e\in A_{br}}|\omega(\eta(e)+(-1)^{\tau(e)}d\eta(\bar{e}))|^{2}$
    $=\frac{1+d^{2}}{2}\sum_{e\in A_{br}}|\eta(e)|^{2}+d\cdot\mathrm{Re}\left[\sum_{e\in A_{br}}(-1)^{\tau(e)}\eta(e)\overline{\eta(\bar{e})}\right]$
    $=\frac{1}{2}\left\{(1+d^{2})\;||\eta||^{2}+2d\cdot\mathrm{Re}(\langle\eta,\sigma\eta\rangle)\right\}$
@@
Remark 7.1.
Let $(G,\rho_{1},\tau_{1})$ and $(G,\rho_{2},\tau_{2})$ be rotation systems which are equivalent to each other. The scattering matrices and the comfortabilities of each rotation system are denoted by $S_{1}$, $S_{2}$ and $\mathcal{E}_{1}$, $\mathcal{E}_{2}$, respectively. By Proposition 6.1, there exists a diagonal unitary matrix $D$ such that $S_{2}=D^{*}S_{1}D$. Then Proposition 7.2 implies that for any inflow $\boldsymbol{\alpha}_{1}$ to the rotation system $(G,\rho_{1},\tau_{1})$, there exists $\boldsymbol{\alpha}_{2}$ such that $\mathcal{E}_{1}$ with the inflow $\boldsymbol{\alpha}_{1}$ coincides with $\mathcal{E}_{2}$ with the inflow $\boldsymbol{\alpha}_{2}$, and the inflow $\boldsymbol{\alpha}_{2}$ is described by $\boldsymbol{\alpha}_{2}=D\boldsymbol{\alpha}_{1}$. Thus $\mathcal{E}_{1}\neq\mathcal{E}_{2}$ with the same inflow in general. However, in the case for a single inflow, that is, $|{\rm supp}(\boldsymbol{\alpha}_{in})|=1$, then $\mathcal{E}_{1}=\mathcal{E}_{2}$, which means that the comfortability with the single inflow is invariant among the equivalent embbedings.
Concerning the observation in Remark 7.1, now let us set the inflow inserting $1$ from a single tail which is chosen uniformly at random, that is, an inflow $\boldsymbol{\alpha}_{in}$ is selected randomly from
$\{\delta_{e}\;|\;e\in A_{br}\}.$
@@
Under Assumption 1, suppose that $\boldsymbol{\alpha}_{in}$ is the random variable distributed uniformly in $\{\delta_{e}\;|\;e\in A_{br}\}$. Then the average of the comfortability is expressed by
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(\overline{a\omega})^{\mathrm{dist}_{f}(e,\bar{e})}\;(1-|a|^{2\mathrm{dist}_{f}(\bar{e},e)})\;\bigg{\}}.$
Here $|f|$ is the length of the facial walk $f$ in $(G,\rho,\tau)$ and $\mathrm{dist}_{f}(e,e^{\prime})$ is the distance from $t(e^{\prime})$ to $o(e^{\prime}$ along the facial walk $f$ in $(G,\rho,\tau)$ for any $e,e^{\prime}\in f$.
@@
By Proposition 7.2, it holds that
$\mathbb{E}[\mathcal{E}]$ $=\frac{1}{|c|^{2}}\mathbb{E}[\;||Q\boldsymbol{\alpha}||^{2}\;]+\frac{1}{2|bc|^{2}}\mathbb{E}[\;||(\sigma+dI)Q\boldsymbol{\alpha}||^{2}\;]$
@@
Now to describe the above RHS more explicitly, let us compute $QQ^{*}$ as follows. Put $P_{f}:=P_{f}(1)$. We remark that $P_{f}(\omega)=\omega P_{f}$ for any $f\in F$ under Assumption 1. Note that $Q=\oplus_{f\in F}Q_{f}$, and for each $f\in F$ with $f=(e_{0},e_{1},\dots,{e_{|f|-1}})$, we have
$\sum_{k=0}^{|f|-1}\tau(e_{k})=0,$ (7.43)
which implies $P_{f}^{|f|}=I_{f}$. Then Theorem 6.1 implies that $Q_{f}$ can be expanded by
@@
Then we have
$Q_{f}Q_{f}^{*}$ $=\frac{|bc|^{2}}{|1-(a\omega)^{|f|}|^{2}}\sum_{k=0}^{|f|-1}\left(\sum_{{\scriptsize\begin{matrix}\ell,m\in\mathbb{Z}_{|f|}\\ \text{ with }m-\ell=k\end{matrix}}}(a\omega)^{m}(\overline{a\omega})^{\ell}\right)P_{f}^{k}.$ (7.44)
Here $w_{k}:=\sum_{\ell,m\in\mathbb{Z}_{|f|},\;m-\ell=k}\;(a\omega)^{m}(\overline{a\omega})^{\ell}$ is reduced to
$w_{k}$ $=\sum_{j=0}^{|f|-1-k}(\overline{a\omega})^{j}(a\omega)^{j+k}+\sum_{j=0}^{k-1}(\overline{a\omega})^{|f|-k+i}(a\omega)^{i}$
$=\frac{1}{|b|^{2}}\left\{(a\omega)^{k}(1-|a|^{2(|f|-k)})+(\overline{a\omega})^{|f|-k}(1-|a|^{2k})\right\}.$
Here we used $|a|^{2}+|b|^{2}=1$. Note that
$(P_{f}^{k})_{e_{\ell},e_{m}}=\begin{cases}(-1)^{\sum_{j=m+1}^{\ell}\tau(e_{j})}&\text{: $\mathrm{dist}_{f}(e_{\ell},e_{m})=k$,}\\ 0&\text{: otherwise.}\end{cases}$
Here “$\sum_{j=p}^{q}\tau(j)$” is defined by the summation of $\tau(j)$’s over the walk along the boundary of the face $f$ from $e_{p}$ to $e_{q}$ for $p,q\in\mathbb{Z}_{|f|}$, that is,
$\sum_{j=p}^{q}\tau(j)=\tau(p)+\tau(p+1)+\cdots+\tau(q).$
This means that if $p>q$, then $\sum_{j=p}^{q}\tau(j)=\tau(p)+\tau(p+1)+\cdots+\tau(q+|f|)$. Then inserting the above expression for $w_{k}$ into (7.44), we obtain
$(Q_{f}Q_{f}^{*})_{e_{\ell},e_{m}}=(-1)^{\sum_{k=m+1}^{\ell}\tau(e_{k})}\frac{|b|^{2}}{|1-(a\omega)^{|f|}|^{2}}\\ \times\left\{(a\omega)^{\ell-m}(1-|a|^{2(|f|-(\ell-m))})+(\overline{a\omega})^{|f|-(\ell-m)}(1-|a|^{2(\ell-m)})\right\},$ (7.45)
for any $\ell,m\in\mathbb{Z}_{|f|}$. Note that “$\ell-m$” and ‘$|f|-(\ell-m)$” keep the modulus of $\mathbb{Z}_{|f|}$ in the above expression. By using (7.45), $\mathrm{tr}(QQ^{*})$ is described by
@@
Let the closed walk on $G^{\tau}$ from $e$ to $\bar{e}$ along the facial walk $f$ be denoted by $(e_{0},e_{1},\dots,e_{s-1})$ with $e_{0}=e$ and $e_{s-1}=\bar{e}$. Then $\mathrm{tr}(QQ^{*}\sigma)$ is described by
$\mathrm{tr}(QQ^{*}\sigma)$ $=\sum_{e}\sum_{e^{\prime}}(QQ^{*})_{e,e^{\prime}}(\sigma)_{e^{\prime},e}=\sum_{e}(QQ^{*})_{\bar{e},e}(-1)^{\tau(e)}$
$=\sum_{f\in F}\sum_{e\in f\cap\bar{f}}(Q_{f}Q_{f}^{*})_{\bar{e},e}(-1)^{\tau(e)}$
@@
×$\quad\times\sum_{e\in f\cap\bar{f}}\left\{(a\omega)^{\mathrm{dist}_{f}(\bar{e},e)}(1-|a|^{2\mathrm{dist}_{f}(e,\bar{e})})+(\overline{a\omega})^{\mathrm{dist}_{f}(e,\bar{e})}(1-|a|^{2\mathrm{dist}_{f}(\bar{e},e)})\right\}(-1)^{\sum_{k=0}^{s-1}\tau(e_{k})}.$ (7.47)
Here in the last equality we used (7.44) and $\mathrm{dist}_{f}(e,\bar{e})=|f|-\mathrm{dist}_{f}(\bar{e},e)$. We remark that $\sum_{k=0}^{s-1}\tau(e)=0$ since $e,\bar{e}$ lie on the same sheet. Combining (7) with (7.46) and (7.47), we obtain the desired conclusion. ∎